Determine whether the given pair is a solution of equations. Remember to use alphabetical ordered variables.
(1,9) 7x-y=-2
3x-7y=-60

Answer by jim_thompson5910(35256)

Take the first equation, 7x-y=-2, plug in x = 1 and y = 9 to get

7x-y=-2

7(1)-(9)=-2


From here, simplify and evaluate. If the end result is true (something like 1=1), then the ordered pair satisfies the first equation.

Rinse and repeat but do so for the second equation. Plug in x = 1 and y = 9 into the second equation and see if a true equation results.


If BOTH equations are satisfied, then (1,9) is indeed a solution.
If either equation (or both) are NOT satisfied, then (1,9) is NOT a solution